### v2

@@ 3.3 Masking
We apply masking to the audio and visual sequences of tokens. The MAE aims to reconstruct the masked tokens from the visible tokens to learn a semantic representation of the data. The masking strategy impacts the performance of downstream tasks. In the original MAE, the masked tokens are randomly drawn with a target ratio of the entire set of tokens, typically 75%. In multimodal scenarios, the basic strategy would be to do the same for each modality, but it has been shown that it is more interesting to implement a coupled masking strategy between the modalities [Bachmann et al., 2022]. This involves randomly drawing a masking proportion $p_{a}\in[0,1]$ for the audio modality and $p_{v}\in[0,1]$ for the visual modality such that $p_{a}+p_{v}=1$. To do this, a Dirichlet distribution is used: $(p_{a},p_{v})\sim Dir(\alpha_{a},\alpha_{v})$ where $\alpha_{a},\alpha_{v}>0$ are the concentration parameters of the distribution. When $\alpha_{a}=\alpha_{v}=1$, the distribution is uniform overall points in its support (the 1-simplex), and $\alpha_{a}\gg\alpha_{b}$ results in a sampling behavior where most of the tokens are taken from the audio modality, and vice versa. These parameters can be used to reduce the dominance of one modality over the other. This strategy allows the reconstruction of missing information from one modality by relying on another. Therefore, the model will leverage the less masked modality for the reconstruction of the more masked one. For this paper, we set $\alpha_{a}=\alpha_{v}=1$.
@@ 3.4 Continuous embedding vectors
The discrete tokens correspond to the indices obtained through the quantization step of the pretrained VQ-VAE encoder. Before being input to the VQ-MAE-AV encoder, these discrete tokens are replaced with trainable continuous embedding vectors taken from an audio codebook in $\mathbb{R}^{k_{a}\times e_{a}}$ and from a visual codebook in $\mathbb{R}^{k_{v}\times e_{v}}$, where $k_{a/v}$ is the number of codes in the codebook and $e_{a/v}$ is the dimension of each code. This is simply achieved by replacing the indices of a discrete token with the corresponding vectors of dimension $e_{a/v}$ in the codebook. After this embedding process, the sequences of discrete tokens ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{Z}^{(n_{t_{a}}\cdot n_{d})\times(t_{a}\cdot d)}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$ in $\mathbb{Z}^{(n_{t_{v}}\cdot n_{h}\cdot n_{w})\times(t_{v}\cdot h\cdot w)}$ are transformed into sequences of continuous tokens ${\color[rgb]{1,0,0}\mathbf{x}_{c}^{(a)}}\in\mathbb{R}^{(n_{t_{a}}\cdot n_{d})\times({t_{a}}\cdot d\cdot e_{a})}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{c}^{(v)}}\in\mathbb{Z}^{(n_{t_{v}}\cdot n_{h}\cdot n_{w})\times({t_{v}}\cdot h\cdot w\cdot e_{v})}$.
@@ 3.5.1 Attention Block
The VQ-MAE-AV encoder and decoder are built with multi-head Attention blocks similar to those used in the Vision Transformer (ViT) [Dosovitskiy et al., 2020]. Each block comprises a multi-head attention layer, normalization layers, and a Multi-layer Perceptron (MLP). These layers are interconnected with residual connections, as depicted in Fig. 2, and they will be used to capture the inter and intra-relationships between audio and visual tokens. This attention block is inspired by the attention layer in the original transformer [Vaswani et al., 2017]. To simplify the reading afterward, we denote the attention block by $\texttt{Attention}(Q,V,K)$, where $Q,V,K$ are the query, value, and key, respectively. The self-attention mechanism uses the same input vector for the query, key, and value vectors. In the case of cross-attention, the query and key are different to enable attention across multiple modalities or inputs.
@@ 3.5.2 Encoders
We propose two fusion strategies of the audio and visual speech data, resulting in two architectures for the VQ-MAE-AV encoder. The first fusion strategy is called self-attention fusion (SAF). As represented in Fig. 2, this fusion consists of a concatenation of token sequences for the two modalities, followed by $L$ self-attention blocks. This concatenation operates on the first dimension of the variables, i.e., we obtain a sequence of $(n_{t_{a}}\cdot n_{d})+(n_{t_{v}}\cdot n_{h}\cdot n_{w})$ tokens after concatenation.
The second fusion strategy is called cross-attention fusion (CAF). As shown in Fig. 2, the sequences of audio and visual tokens are used separately as the queries of two separate attention blocks, which share the same keys and values corresponding to the concatenation of the modalities. These two cross-attention blocks are then followed by a stack of $L$ self-attention blocks.
@@ 3.5.3 Global tokens
In addition to the token-wise representations ${\color[rgb]{1,0,0}\mathbf{z}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$, we learn two sequence-wise global tokens, denoted by ${\color[rgb]{1,0,0}\mathbf{w}^{(a)}}\in\mathbb{R}^{1\times(t_{a}\cdot d\cdot e_{a})}$ and ${\color[rgb]{0,0,1}\mathbf{w}^{(v)}}\in\mathbb{R}^{1\times({t_{v}}\cdot h\cdot w\cdot e_{v})}$, which can be thought of as similar to [CLS] tokens [He et al., 2022]. These global modality-specific tokens are introduced to aggregate the spectro-temporal and spatio-temporal information in the two modalities, which can be useful for downstream tasks involving predictions at the sequence level, such as audiovisual SER. The global tokens are computed with attention pooling as follows:
@@
where $Q_{(a)}=\mathcal{Z}^{(a)}\in\mathbb{R}^{1\times(t_{a}\cdot d\cdot e_{a})}$ and $Q_{(v)}=\mathcal{Z}^{(v)}\in\mathbb{R}^{1\times({t_{v}}\cdot h\cdot w\cdot e_{v})}$ represent respectively trainable audio and visual tokens, as proposed in Touvron et al. [2021], $V_{(a)}=K_{(a)}={\color[rgb]{1,0,0}\mathbf{z}^{(a)}}$, and $V_{(v)}=K_{(v)}={\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$.
@@ 3.5.4 Decoders
The token-wise representation obtained from the encoder is combined with mask tokens and fed to the VQ-MAE-AV decoder along with additional position embeddings, as denoted by ${\color[rgb]{1,0,0}\tilde{\mathbf{z}}^{(a)}}$ and ${\color[rgb]{0,0,1}\tilde{\mathbf{z}}^{(v)}}$ and illustrated in Fig. 2. The mask tokens actually correspond to one single trainable vector as proposed in the original MAE [He et al., 2022]. Similarly as for the encoder, the audio and visual inputs of the VQ-MAE-AV decoder can be fused using either self-attention fusion or cross-attention fusion.
@@
A linear layer is added at the end of the decoder, which maps to the size of the VQ-VAE codebooks. The output of this linear layer corresponds to the logits of the discrete tokens. After applying a softmax operation, we obtain reconstructions ${\color[rgb]{1,0,0}\mathbf{\hat{x}}^{(a)}_{q}}$ and ${\color[rgb]{0,0,1}\mathbf{\hat{x}}^{(v)}_{q}}$ of the indices ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$ that were provided by the VQ-VAE-audio and VQ-VAE-visual encoders, respectively.
3.6 VQ-MAE-AV loss functions
3.6.1 Generative loss function
To train the VQ-MAE-AV model, we minimize the cross-entropy loss applied only to the masked discrete tokens:
$\mathcal{L}_{rec}=\texttt{cross-entropy}\Biggl{(}{\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\Big{(}\Omega_{\mathcal{M}}^{(a)}\Big{)},~{}{\color[rgb]{1,0,0}\mathbf{\hat{x}}^{(a)}_{q}}\Big{(}\Omega_{\mathcal{M}}^{(a)}\Big{)}\Biggl{)}$
[…]
where $\mathbf{x}\left(\Omega_{\mathcal{M}}^{(\cdot)}\right)$ denotes the set of masked tokens in $\mathbf{x}$. Another benefit of manipulating discrete representations for multimodal inputs is the homogeneity of the losses, which does not require balancing the losses between the two modalities.
3.6.2 Contrastive loss function
Building upon the approaches presented in Alayrac et al. [2020], Akbari et al. [2021], the global tokens are learned using noise contrastive estimation. This approach enhances the alignment of audiovisual speech pairs by grouping together embeddings that belong to the same time sequence and separating them from those that do not correspond to the same sequence. This approach involves minimizing the loss function $\mathcal{L}_{NCE}({\color[rgb]{1,0,0}\mathbf{w}^{(a)}},{\color[rgb]{0,0,1}\mathbf{w}^{(v)}})$, which is defined by:
$\mathcal{L}_{NCE}(\mathbf{u},\mathbf{v})=$
$-\log\left(\frac{\exp\Bigl{(}\mathbf{u}^{\top}\mathbf{v}/\tau\Bigl{)}}{\exp\Bigl{(}\mathbf{u}^{\top}\mathbf{v}/\tau\Bigl{)}+\sum\limits_{(\mathbf{u}^{\prime},~{}\mathbf{v}^{\prime})\in\mathcal{N}}\exp\Bigl{(}\mathbf{u}^{\prime\top}\mathbf{v}^{\prime}/\tau\Bigl{)}}\right).$ (4)
@@ 3.6.2 Contrastive loss function
3.6.3 Final loss function
The final loss function corresponds to the sum of the generative loss function in equation (3.6.1) and of the contrastive loss function in equation (3.6.2).
(a) Qualitative unmasking results for the audio modality. The spectrogram highlighted in the red box represents the original spectrogram. The two spectrograms on the top right represent the spectrograms masked at 50% and 80%, respectively. The reconstructions using VQ-MAE-AV can be seen directly below these masked spectrograms.
(b) Qualitative unmasking results for the visual modality. The first sequence shows the original video, followed by the next two sequences representing the masked video with a ratio of 50% and its reconstruction using VQ-MAE-AV. The last two sequences represent the masked video with a ratio of 80% and its reconstruction using VQ-MAE-AV.
Figure 3: Quantitative results of the audio reconstruction (a) and visual reconstruction (b) using the VQ-MAE-AV model.
3.7 Fine-tuning for audiovisual SER
After pre-training VQ-MAE-AV on token unmasking, the model is fine-tuned for audiovisual SER. We propose three different approaches for fine-tuning the model. The first two approaches rely on a pooling operation (of the encoder outputs ${\color[rgb]{1,0,0}\mathbf{z}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$) to extract the global audio and visual tokens, which is either a simple mean pooling or the attention pooling discussed in Section 3.5.3. The extracted tokens are then concatenated and passed through a single linear layer, followed by an $argmax$ operation to classify the emotion. The third proposed approach, referred to as Query2Emo and inspired by [Liu et al., 2021], involves cross-attention between all audio and visual tokens (concatenation of ${\color[rgb]{1,0,0}\mathbf{z}^{(a)}},~{}{\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$) as key and value, and the emotion classes represented by trainable embeddings as the query. Query2Emo has a single attention block for both the encoder and decoder. The extracted embeddings are then concatenated and passed through a single linear layer. We adopt the asymmetric loss [Ridnik et al., 2021] for all these approaches between the predicted emotion and the ground-truth emotion.

### v3

@@ 3.3 Masking
The masking strategy employed to train the MAE will impact the performance of downstream tasks He et al. (2022). In the original MAE, the masked tokens are chosen randomly given a target masking ratio, typically 75%. The MAE is then trained to reconstruct the masked tokens from the visible ones, and in doing so it learns a representation that can be used for downstream tasks. In a multimodal scenario, the straightforward approach would be to apply this same masking strategy independently on each modality. However, it has been shown that it is more effective to implement a coupled masking strategy between the modalities (Bachmann et al., 2022). Therefore, to the train the VQ-MAE-AV model, the masking ratio for the audio and visual modalities is drawn randomly according to a uniform distribution on the 1-simplex. This means that if $p\times 100~{}\%$ of the tokens are masked in one modality, then $(1-p)\times 100~{}\%$ of the tokens are masked in the other modality, where $p$ is distributed uniformly between 0 and 1. This strategy allows the reconstruction of missing information from one modality by relying on the other.
Fig. 3: Overview of the three emotion recognition models trained on top of the VQ-MAE-AV encoder.
@@ 3.4 Continuous embedding vectors
The discrete tokens correspond to the indices obtained through the quantization step of the pretrained VQ-VAE encoder. Before being input into the VQ-MAE-AV encoder, these discrete tokens are replaced with trainable continuous embedding vectors taken from an audio codebook in $\mathbb{R}^{k_{a}\times e_{a}}$ and from a visual codebook in $\mathbb{R}^{k_{v}\times e_{v}}$, where $k_{a/v}$ is the number of codes in the codebook and $e_{a/v}$ is the dimension of each code. This is simply achieved by replacing the indices of a discrete token with the corresponding vectors of dimension $e_{a/v}$ in the codebook. After this embedding process, the sequences of discrete tokens ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{N}^{(n_{t_{a}}\cdot n_{d})\times(t_{a}\cdot d)}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$ in $\mathbb{N}^{(n_{t_{v}}\cdot n_{h}\cdot n_{w})\times(t_{v}\cdot h\cdot w)}$ are transformed into sequences of continuous tokens ${\color[rgb]{1,0,0}\mathbf{x}_{c}^{(a)}}\in\mathbb{R}^{(n_{t_{a}}\cdot n_{d})\times({t_{a}}\cdot d\cdot e_{a})}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{c}^{(v)}}\in\mathbb{R}^{(n_{t_{v}}\cdot n_{h}\cdot n_{w})\times({t_{v}}\cdot h\cdot w\cdot e_{v})}$. The dimensions $e_{a}$ and $e_{v}$ of the audio and visual codes are chosen such that the continuous tokens in both modalities have the same dimension $E={t_{a}}\cdot d\cdot e_{a}={t_{v}}\cdot h\cdot w\cdot e_{v}$, which is necessary for concatenation in the VQ-MAE-AV encoder (see the next subsection). Therefore, we have ${\color[rgb]{1,0,0}\mathbf{x}_{c}^{(a)}}\in\mathbb{R}^{(n_{t_{a}}\cdot n_{d})\times E}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{c}^{(v)}}\in\mathbb{R}^{(n_{t_{v}}\cdot n_{h}\cdot n_{w})\times E}$.
@@ 3.5.1 Attention Block
The VQ-MAE-AV encoder and decoder are built with multi-head Attention blocks similar to those used in the Vision Transformer (ViT) (Dosovitskiy et al., 2020). Each block comprises a multi-head attention layer, normalization layers, and a Multi-layer Perceptron (MLP). These layers are interconnected with residual connections, as depicted in Fig. 2, and they will be used to capture the inter and intra-relationships between audio and visual tokens. This attention block is inspired by the attention layer in the original transformer (Vaswani et al., 2017). To simplify the reading afterward, we denote the attention block by $\texttt{Attention}(Q,V,K)$, where $Q,V,K$ are the query, value, and key, respectively. The self-attention mechanism uses the same input vector for the query, key, and value vectors. In the case of cross-attention, the query and key are different to enable attention across multiple modalities or inputs.
@@ 3.5.2 Encoders
We propose two fusion strategies of the audio and visual speech data, resulting in two architectures for the VQ-MAE-AV encoder. The first fusion strategy is called self-attention fusion. As represented in Fig. 2, this fusion consists of a concatenation of token sequences for the two modalities, followed by $L$ self-attention blocks. This concatenation operates on the first dimension of the variables, i.e., we obtain a sequence of $(n_{t_{a}}\cdot n_{d})+(n_{t_{v}}\cdot n_{h}\cdot n_{w})$ tokens after concatenation.
The second fusion strategy is called cross-attention fusion. As shown in Fig. 2, the sequences of audio and visual tokens are used separately as the queries of two separate attention blocks, which share the same keys and values corresponding to the concatenation of the modalities. These two cross-attention blocks are then followed by a stack of $L$ self-attention blocks.
@@ 3.5.3 Global tokens
In addition to the token-wise representations ${\color[rgb]{1,0,0}\mathbf{z}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$, we learn two sequence-wise global tokens, denoted by ${\color[rgb]{1,0,0}\mathbf{w}^{(a)}}\in\mathbb{R}^{1\times E}$ and ${\color[rgb]{0,0,1}\mathbf{w}^{(v)}}\in\mathbb{R}^{1\times E}$, which can be thought of as similar to [CLS] tokens (He et al., 2022). These global modality-specific tokens are introduced to aggregate the spectro-temporal and spatio-temporal information in the two modalities, which can be useful for downstream tasks involving predictions at the sequence level, such as audiovisual SER. The global tokens are computed with attention pooling as follows:
@@
where $Q_{(a)}\in\mathbb{R}^{1\times E}$ and $Q_{(v)}\in\mathbb{R}^{1\times E}$ represent respectively trainable audio and visual tokens, as proposed in Touvron et al. (2021), $V_{(a)}=K_{(a)}={\color[rgb]{1,0,0}\mathbf{z}^{(a)}}$, and $V_{(v)}=K_{(v)}={\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$.
@@ 3.5.4 Decoders
The token-wise representation obtained from the encoder is combined with mask tokens and fed to the VQ-MAE-AV decoder along with additional position embeddings, as denoted by ${\color[rgb]{1,0,0}\tilde{\mathbf{z}}^{(a)}}$ and ${\color[rgb]{0,0,1}\tilde{\mathbf{z}}^{(v)}}$ and illustrated in Fig. 2. The mask tokens actually correspond to one single trainable vector as proposed in the original MAE (He et al., 2022). Similarly as for the encoder, the audio and visual inputs of the VQ-MAE-AV decoder can be fused using either self-attention fusion or cross-attention fusion.
@@
A linear layer is added at the end of the decoder, which maps to the size of the VQ-VAE codebooks. The output of this linear layer corresponds to the logits of the discrete tokens. After applying a argmax operation, we obtain reconstructions ${\color[rgb]{1,0,0}\mathbf{\hat{x}}^{(a)}_{q}}$ and ${\color[rgb]{0,0,1}\mathbf{\hat{x}}^{(v)}_{q}}$ of the indices ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$ that were provided by the VQ-VAE-A and VQ-VAE-V encoders, respectively.
3.6 Loss functions
The VQ-MAE-AV model is trained in a self-supervised manner to minimize a generative loss $\mathcal{L}_{gen}$ between the reconstructed and original tokens and a contrastive loss $\mathcal{L}_{NCE}$ between the audio and visual global tokens.
Due to their discrete nature, the cross-entropy (CE) is naturally used to measure the reconstruction quality of the audiovisual masked tokens:
$\mathcal{L}_{gen}$ $\Big{(}{\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}},{\color[rgb]{1,0,0}\mathbf{\hat{x}}^{(a)}_{q}},{\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}},{\color[rgb]{0,0,1}\mathbf{\hat{x}}^{(v)}_{q}},\Omega^{(a)},\Omega^{(v)}\Big{)}=$
$\texttt{CE}\Biggl{(}{\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\Big{(}\Omega^{(a)}\Big{)},~{}{\color[rgb]{1,0,0}\mathbf{\hat{x}}^{(a)}_{q}}\Big{(}\Omega^{(a)}\Big{)}\Biggl{)}\,+\,\texttt{CE}\Biggl{(}{\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\Big{(}\Omega^{(v)}\Big{)},~{}{\color[rgb]{0,0,1}\mathbf{\hat{x}}^{(v)}_{q}}\Big{(}\Omega^{(v)}\Big{)}\Biggl{)},$ (3)
where $\mathbf{x}\left(\Omega^{(\cdot)}\right)$ denotes the set of masked tokens in $\mathbf{x}$. As can be seen, another benefit of manipulating discrete representations for multimodal inputs is the homogeneity of the losses, which does not require balancing the losses between the two modalities.
Building upon the approaches presented in Alayrac et al. (2020); Akbari et al. (2021), the global tokens are learned using noise contrastive estimation. This approach enhances the alignment of audiovisual speech pairs by grouping together embeddings that belong to the same time sequence and separating them from those that do not correspond to the same sequence. It involves minimizing the loss function $\mathcal{L}_{NCE}({\color[rgb]{1,0,0}\mathbf{w}^{(a)}},{\color[rgb]{0,0,1}\mathbf{w}^{(v)}})$, which is defined by:
$\mathcal{L}_{NCE}(\mathbf{u},\mathbf{v})=-\log\left(\frac{\exp\Bigl{(}\mathbf{u}^{\top}\mathbf{v}/\tau\Bigl{)}}{\exp\Bigl{(}\mathbf{u}^{\top}\mathbf{v}/\tau\Bigl{)}+\sum\limits_{(\mathbf{u}^{\prime},~{}\mathbf{v}^{\prime})\in\mathcal{N}}\exp\Bigl{(}\mathbf{u}^{\prime\top}\mathbf{v}^{\prime}/\tau\Bigl{)}}\right).$ (4)
@@ 3.6 Loss functions
The overall loss function used to train VQ-MAE-AV is simply the sum of the generative and contrastive loss functions in (3.6) and (4).
3.7 Emotion recognition
After self-supervised learning on a large-scale unlabeled dataset, the VQ-MAE-AV model is used for emotion recognition. The task is to predict the emotion class of an input audiovisual speech sequence. To address this task, we need to introduce a small emotion recognition model for predicting the emotion category from the audiovisual speech representation provided by the pre-trained VQ-MAE-AV encoder. In this work, we propose and investigate three different emotion recognition models, which are described in the remainder of this section and illustrated in Fig. 3. All emotion recognition models are trained for a supervised classification task using the asymmetric loss of Ridnik et al. (2021). Along with the training of the emotion recognition model, we also fine-tune the VQ-MAE-AV encoder.
3.7.1 Emotion recognition model based on attention pooling
The first approach to performing emotion recognition from the proposed VQ-MAE-AV model is to use the sequence-wise global tokens ${\color[rgb]{1,0,0}\mathbf{w}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{w}^{(v)}}$, which were computed using attention pooling of the token-wise representations ${\color[rgb]{1,0,0}\mathbf{z}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$, as described in Section 3.5.3. In this approach, the two global tokens are concatenated and passed through a single linear layer followed by a softmax operation to predict the emotion class probabilities.
3.7.2 Emotion recognition model based on mean pooling
The second approach for emotion recognition replaces the previous attention pooling strategy with a more naive one, where the token-wise representations ${\color[rgb]{1,0,0}\mathbf{z}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$ are aggregated temporally using a simple parameter-free mean pooling operation. The resulting vectors are then concatenated and passed through a single linear layer, as before.
3.7.3 Emotion recognition model based on Query2Emo
Finally, we propose a third emotion recognition approach, referred to as Query2Emo and inspired by (Liu et al., 2021). As illustrated in Fig. 3, Query2Emo involves cross-attention between all audio and visual tokens (concatenation of ${\color[rgb]{1,0,0}\mathbf{z}^{(a)}},~{}{\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$) as key and value, and the emotion classes represented by trainable embeddings as the query:
$\mathbf{w}_{emo}=\texttt{Attention}\left(Q_{emo},\mathbf{z},\mathbf{z}\right),$ (5)
where $\mathbf{z}$ denotes the concatenation of the two sequences ${\color[rgb]{1,0,0}\mathbf{z}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$, and $Q_{emo}\in\mathbb{R}^{K_{emo}\times E}$ corresponds to $K_{emo}$ trainable tokens of dimension $E$, with $K_{emo}$ the number of emotion classes. This cross-attention aims to learn the relationship between the emotion representation $Q_{emo}$ and the audiovisual tokens for the SER task. Query2Emo consists of two attention blocks as illustrated in Fig. 3. The model provides at the output $K_{emo}$ vectors of dimension $E$ (i.e., $\mathbf{w}_{emo}\in\mathbb{R}^{K_{emo}\times E})$, which are then concatenated and passed through a single linear layer, as for the two previous emotion recognition models.